# Experiential Explanations for Reinforcement Learning

### v4

@@ 7 Conclusions
Explaining reinforcement learning to non-AI expert audiences is challenging. Actionable local explanations should reference future anticipated interactions with the environment. Our Experiential Explanations technique attempts to capture the “experiences” of the agent in terms of reward influences that can be given to the user in the form of explanations.
Quantitative and qualitative studies with non-AI expert study participants show that Experiential Explanations help users better understand the agent and predict its behaviors. They also show that Experiential Explanations are preferred and found more useful than common alternatives used in many other systems.
Experiential explanations provide context to the RL agent’s behaviors without changing the agent architecture. This is significant because any core RL algorithm can be used to train the influence predictors, as we demonstrated through our studies using both Deep $Q$-Networks and PPO.
The experimental results suggest that Experiential Explanations may provide for more actionable ways for non-AI expert users to understand the behavior of reinforcement learning agents.Explainable RL will be a crucial component as reinforcement learning algorithms begin to drive agents and robots that will cross paths with humans in everyday life.
@@ Appendix D Experiment 1: Evaluation Survey
    In this section, we show a preview of the evaluation survey. Figure 10 shows an example of a prediction task. Then after the participants submit their predictions, they see the correct answer as shown in Figure 11. Afterward, they see one of the three explanation types, Experiential Explanation (Figure 12), Heatmap Explanation (Figure 13), Q-value explanation (Figure 14), or they see no explanation at all. Then they repeat this process two times. After the participants finish the prediction task, they proceed to the reflection survey, where they evaluate their experience with the shown explanations as in figure 15.
@@
    Appendix E Satisfaction Survey Analysis
    Figure 16 shows the results of the satisfaction survey questions shown in 15. The discussion and analysis of the figure 16 results are in the main paper.
    Figure 16: The satisfaction survey results for each explanation quality dimensions grouped by explanation type

### v5

@@ 8 Conclusions
Explaining reinforcement learning to non-AI expert audiences is challenging. Actionable local explanations should reference future anticipated interactions with respect to the environment. Our Experiential Explanations technique attempts to capture the “experiences” of the agent in terms of reward influences that can be relayed to the user in the form of explanations that help them construct an appropriate mental model of how the agent interacts with the environment. Experiential Explanations provide context to the RL agent’s behaviors without changing the agent architecture while still providing a high degree of faithfulness. This is significant because any core RL algorithm can be used to train the influence predictors, as we demonstrated through our studies using both Deep $Q$-Networks and PPO. Quantitative and qualitative studies with non-AI expert study participants show that Experiential Explanations help users better understand the agent and predict its behaviors in partially-observable long-horizon discrete planning problems with sparse rewards. They also show that Experiential Explanations are preferred and found more useful than the compared agent explanation baselines. The experimental results suggest that Experiential Explanations may provide for more actionable ways for non-AI expert users to understand the behavior of reinforcement learning agents.
Appendix A Implementation details of the Minigrid Agent
Figure 14: The architecture of the agent A2C model
Figure 14 shows the architecture of our agent. The state observation is encoded using a convolutional neural network (CNN), then passed to the actor-critic network heads composed of two shallow MLPs with one hidden layer each. The Minigrid agent was implemented using the torch-ac implementation presented in the official Minigrid documentation.  [^6] https://github.com/lcswillems/rl-starter-files/
We train our agents with the following hyperparameters:
processes $=20$
frames $=1221120$
batch size $=256$
discount factor $=0.99$
learning rate $=3.5\times 10^{-4}$
PPO hyperparameters:
epochs $=10$
lambda coefficient in GAE $=0.95$
entropy term coefficient $=0.01$
value loss term coefficient $=0.5$
maximum norm of the gradient $=0.5$
optimizer $=Adam$
optimizer epsilon $=1\times 10^{-8}$
clipping epsilon for PPO $=0.2$
@@ Appendix D Experiment 1: Evaluation Survey
The Human Evaluation Study in Minigrid is made up of three parts. The first part is a prediction task, as shown in the example in Figure 16. We presented three episodes to each participant, as shown in Figure 3. During each episode, the agent—the red triangle—must navigate to the green square. Some colored shapes are associated with negative rewards, but the participant is not told which ones. After each episode, the participant is asked to predict whether the agent will go up or down. The participant has no information at the beginning of the first episode to make an informed choice, and the participant choice is random—we do not evaluate on this choice. Further, upon the participant’s choice, we assign which object is dangerous such that the participant is always wrong. This established the conditions for which the participant must now learn to correct their understanding of the agent’s interactions with the environment. We use that assignment of dangerous object consistently throughout the second and third episodes. The participant tries to predict the agent’s choice after the second and third episodes (without the object assignment switching); we expect to see the participants in the experimental condition to have higher accuracy rates than in the baseline conditions.
After each episode, participants fill out a brief text entry question explaining the rationale behind their choice. Then after the participants submit their predictions and reasoning, they see the correct answer as shown in Figure 17. Afterward, they see one of the three types of explanations, Experiential Explanation (Figure 18), Heatmap Explanation (Figure 19), Q-value explanation (Figure 20), or they see no explanation at all. Then they repeat this process two times.
After the participants finish the prediction task, they proceed to the second part of the study, a satisfaction survey, consisting of five-point Likert scale questions asking the user to rate understandability, satisfaction, amount of details, completeness, usefulness, precision, and trust for the explanations they saw. It also had an open-ended response question asking how the explanation helped or did not help them understand the agent. This survey is shown in figure 21.
@@
Appendix E Experiment 1: Satisfaction Survey Analysis
Figure 22 shows the results of the satisfaction survey questions shown in Figure 21. The discussion and analysis of the Figure 22 results are in the main paper.
Figure 22: Experiment 1 satisfaction survey results for each explanation quality dimensions grouped by explanation type
@@ Appendix F Crafter Re-skin mappings
Appendix G Experiment 2: Evaluation Survey
The Human Evaluation Study in Crafter has three tasks, a prediction task and an ordering task, and a satisfaction survey. We picked four random decision points from the agent’s trajectory, each paired with a counterfactual trajectory that the agent can take by performing a counterfactual action at this decision point. We presented each participant with one decision point and asked them to predict the agent’s action as shown in Figure 23. After this first interaction, they are shown their assigned explanation. The explanations are one of the three baselines specified in Section 6.3. An agent value explanation (Figure 24), the local experiential explanation (Figure 25), or they do not see an explanation.
Then after seeing the explanation, we asked them to predict and reason about the agent’s action at a future decision point displayed in Figure 26. Lastly, the participants engage in an additional task in order of the influences by their importance to the agent when it considers performing the counterfactual action or the actual action as shown in Figure 27. Both tasks are followed by a text entry question asking them to reason about their predictions.
After participants complete both tasks, they proceed to the last part of the study, a satisfaction survey identical to the one we used in experiment 1 as shown in Figure 21. consisting of five-point Likert scale questions asking the user to rate understandability, satisfaction, amount of details, completeness, usefulness, precision, and trust for the explanations they saw. It also had an open-ended response question asking how the explanation helped or did not help them understand the agent. The results of the survey for Experiment 2 is shown in Figure 22.
Figure 23: The first decision point the participants encounter during experiment 2.
Figure 24: An example of the agent value explanation baseline.
Figure 25: An example of the local experiential explanation.
Figure 26: The second decision point the participants encounter during experiment 2. Which is used for the prediction task and followed by the reasoning question.
Figure 27: The ordering task which asks the participants to order the influences by their importance to the agent at the decision point.
Figure 28: Experiment 2 satisfaction survey results for each explanation quality dimensions grouped by explanation type